OpenStudy (anonymous):

The graph of one particular link includes the two points (5,8) and (10,6). Find the slope of the line. 2/5 -2/5 5/2 -5/2

OpenStudy (anonymous):

@inkyvoyd

OpenStudy (inkyvoyd):

use slope formula delta y/ delta x=slope

OpenStudy (anonymous):

@iPwnBunnies

OpenStudy (ipwnbunnies):

What Inky said. The slope of a line between two points is: \[\frac{Y2-Y1}{X2-X1}\]

OpenStudy (anonymous):

tell me wat it is i don't have much time the device is about to explode

OpenStudy (ipwnbunnies):

Breh. (6-8)/(10-5) = -2/5

OpenStudy (anonymous):

ya i new it was b